OpenStudy (anonymous):

There are six chairs in a row. Person A picks one of the chairs at random. Then Person B picks a chair at random from the five chairs that remain. Find the chance that they pick chairs that are next to each other.

jimthompson5910 (jim_thompson5910):

Let's label the 6 chairs as 6 slots The chair on the very left is slot 1. The chair furthest on the right is slot 6. ----------------------------------------------------------------- Case 1) A chooses slot 1 P(A chooses slot 1) = 1/6 P(B chooses slot 2) = 1/5 Multiply them out to get (1/6)*(1/5) = 1/30 ----------------------------------------------------------------- Case 2) A chooses slot 6 P(A chooses slot 6) = 1/6 P(B chooses slot 5) = 1/5 Multiply them out to get (1/6)*(1/5) = 1/30 ----------------------------------------------------------------- Case 3) A chooses slots 2, 3, 4, or 5 P(A chooses slots 2, 3, 4, or 5) = 4/6 = 2/3 P(B chooses a slot next to A) = 2/5 ... Note: there are 2 chairs on either side of slot 2, 3, 4, or 5 Multiply them out to get (2/3)*(2/5) = 4/15

OpenStudy (anonymous):

120?

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

chance...

jimthompson5910 (jim_thompson5910):

Now add up the probabilities found in cases 1 through 3 1/30 + 1/30 + 4/15 1/30 + 1/30 + 8/30 10/30 1/3

OpenStudy (anonymous):

Yes that is correct